A student multiplied 7236 by 65 instead of multiplying by 56. By how much was his answer greater than the correct answer?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Difference between 65 and 56 i.e (65–56) = 9
The difference between the correct and incorrect answer = 7236×9 = 65124 Hence, by 65124, the answer was greater than the correct answer

Thanks